A bi-concave lens has a radius of curvature of one side equal to 10.5cm and on the other side equal to 15.0cm.If the lens is made of crown glass(n=1.52), what is the focal length of the lens?

1
Expert's answer
2021-12-20T10:31:54-0500

"R_1=10.5cm"

"R_2=15cm"

"n=1.52"

"\frac{1}{F}=(n-1)(\frac{1}{R_2}+\frac{1}{R_1}) \to F = \large\frac{R_2*R_1}{(n-1)(R_1+R_2)}=\frac{10.5*15}{0.52*25.5}=" "11.88"